Total Population 2006 All Deaths 2006 Total Accidents 2006 Deaths from Accidents 2006 Death Rate (per...

Total

Population

2006

All Deaths

2006

Total Accidents 2006

Deaths from Accidents

2006

Death Rate

(per 100,000)

Accident-Specific Death Rates

(per 100,000)

Accident Case-Fatality Rate

PMR (%)

5-14

39,976,619

6,149

109,654

2,258

1. Calculate age specific death rate, accident specific death rate and proportionate mortality ratios (PMR). Show your work and calculations for each type.

a. Age-specific death rate = # of deaths per age stratum / Population per age stratum

b. Accident-specific death rate = # of deaths from accidents per age stratum / Population per age stratum

c. Accident case-fatality rate = # of deaths from accidents per age stratum / Total # of accidents per age stratum

d. Mortality Rate = # of deaths / Population at Risk

e. Proportionate Mortality Ratio = # of deaths from accidents per age stratum / Total # of deaths per age stratum
Answer #1

Assuming that given population belongs to a particular age group, Different rates and ratios can be calculated as mentioned below:

1. Age specific death rate = No. of deaths per age stratum / Population per age stratum

  • No. of deaths = 6149
  • Total population = 39,976,619

Age specific death rate = (6149/39,976,619) x 100000

Age specific death rate = 15.382 per 100000 population

2. Accident-specific death rate = No. of deaths from accidents per age stratum / Population per age stratum

  • No. of deaths from accidents = 2258
  • Total population = 39,976,619

Accident-specific death rate = (2258/39,976,619) x 100000

Accident-specific death rate = 5.649 per 100000 population

3. Accident case-fatality rate = No. of deaths from accidents per age stratum / Total no. of accidents per age stratum

  • No. of deaths from accidents = 2258
  • Total no. of accidents = 109,654

Accident case-fatality rate = (2258/109654) x 100000

Accident case-fatality rate = 2059.20 per 100000 population or 20.59 per 1000 population or 2.059 %

4. Mortality Rate = No. of deaths / Population at Risk

  • No. of deaths = 6149
  • Total population = 39,976,619

Mortality Rate = (6149/39,976,619) x 1000 (Mortality rate is expressed per 1000 population)

Mortality Rate = 0.153 per 1000 population or 0.015% (15.382 per 100000 population)

5. Proportionate Mortality Ratio = No. of deaths from accidents per age stratum / Total no. of deaths per age stratum

  • No. of deaths from accidents = 2258
  • Total no. of deaths = 6149

Proportionate Mortality Ratio = (2258/6149) x 100

Proportionate Mortality Ratio = 36.72 %